OpenStudy (anonymous):

There are 2 leaves along 3 in. of an ivy vine. There are 14 leaves along 15 in. of the same vine. Which equation models the number of leaves y along x in. of vine?

OpenStudy (armyrangers):

What Are the choses

OpenStudy (anonymous):

y=-1+x ,y=x-1,y=-x-1,y=x+1

OpenStudy (armyrangers):

Ok so what do you think it is

OpenStudy (anonymous):

-x-1

OpenStudy (armyrangers):

No cause that would make it more leaves then inches

OpenStudy (anonymous):

x+1

OpenStudy (armyrangers):

No so it says that x=inches and y= leaves so lets try it like this plugin 2 for y and 3 for X and then tell me what one leaves you with 2=2

OpenStudy (anonymous):

x-1 the answer

OpenStudy (armyrangers):

Yes that is right good job

OpenStudy (anonymous):

try this Identify the function in which y varies directly with x.